OpenStudy (anonymous):

Perform the indicated operation and simplify. Come to the problem so it can be displayed as an equation:) Thanks and please show work!

OpenStudy (anonymous):

\[((x^2-2x-24)\div(x^2+3x -10)) \times ((3x^2-6x)\div(x^3+4x^2))\]

OpenStudy (dape):

\[ \frac{x^2-2x-24}{x^2+3x-10}*\frac{3x^2-6x}{x^3+4x^2} = \frac{(x-6)(x+4)}{(x-2)(x+5)}*\frac{3x(x-2)}{x^2(x+4)} = \\ = \frac{x-2}{x-2}*\frac{x+4}{x+4}*\frac{x}{x}*\frac{3(x-6)}{x(x+5)} = \frac{3(x-6)}{x(x+5)} = \frac{3x-18}{x^2+5x} \]

OpenStudy (anonymous):

Thanks do you think you can help me with a few more?

OpenStudy (dape):

If they are similar just follow the same basic procedure. See if you can factor the numerator and denominator, then remove the factors that cancel each other (i.e. (x+2)/(x+2)).

OpenStudy (anonymous):

\[((x^2-10x+16)\div (x^2-1)) \times (x-1)\div(1))(\]

OpenStudy (anonymous):

oh sorry about the last parenthesis. My answer was (x-8)(x-2)/(x+1)

OpenStudy (dape):

That is correct, you got it!